For a sample of 50 firms taken from a particular industry the mean number of employees per firm is 420.4 with a sample standard deviation of 55.7. There is a total of 380 firm in this industry.


(i) Determine the standard error of the mean to be used in conjunction with estimating the population mean by a confidence interval.


1
Expert's answer
2021-07-27T15:25:21-0400

For a sample of 50 firms taken from a particular industry the mean number of employees per firm is 420.4 with a sample standard deviation of 55.7.

"n=50, s=55.7"


"se=\dfrac{s}{\sqrt{n}}=\dfrac{55.7}{\sqrt{50}}\approx7.8772"